# ViGoR: Improving Visual Grounding of Large Vision Language Models with Fine-Grained Reward Modeling

### v1

@@ Abstract
By combining natural language understanding and the generation capabilities and breadth of knowledge of large language models with image perception, recent large vision language models (LVLMs) have shown unprecedented reasoning capabilities in the real world. However, the generated text often suffers from inaccurate grounding in the visual input, resulting in errors such as hallucinating nonexistent scene elements, missing significant parts of the scene, and inferring incorrect attributes and relationships between objects. To address these issues, we introduce a novel framework, ViGoR (Visual Grounding Through Fine-Grained Reward Modeling) that utilizes fine-grained reward modeling to significantly enhance the visual grounding of LVLMs over pre-trained baselines. This improvement is efficiently achieved using much cheaper human evaluations instead of full supervisions, as well as automated methods. We show the effectiveness of our approach through numerous metrics on several benchmarks. Additionally, we construct a comprehensive and challenging dataset specifically designed to validate the visual grounding capabilities of LVLMs. Finally, we plan to release our human annotation comprising approximately 16,000 images and generated text pairs with fine-grained evaluations to contribute to related research in the community.
@@ 1 Introduction
Figure 1: An illustration of inaccurate visual grounding in Large Vision Language Models. BLIP-2 failed to provide detailed image descriptions while LLaVA offers more detail but lacks accuracy and frequently results in hallucinations. In contrast, our model preserves the logical reasoning and creativity of large language models while exhibiting significantly better accuracy and detail.
@@
Large language models (LLMs) have garnered intense interest throughout the research and academic communities. Typically, these models are pre-trained with tremendous amounts of automatically aggregated text data and are further fine-tuned with specific user examples or evaluation feedback to enable the models to follow human-provided prompts and retrieve useful relevant information or solve logical problems. Recently, numerous techniques have augmented these breakthroughs with the ability to understand visual information by further integrating image features into the prompt encoding process. Although these works have successfully aligned image features into the large language model domain, they still exhibit severe problems. While a strong contextual grounding of language-only models can be learned from the enormous corpus of text data, training data for multimodal language/vision models is much more limited and flawed, while the complexity of the task is arguably higher. Large, automatically compiled datasets, such as the LAION-5B dataset [27], tend to feature only simple images with very short text descriptions. Training with such data often leads large vision language models to fail in capturing the essential details of the image, returning only a short and coarse description (see BLIP-2 [16] output in Figure 1). Moreover, techniques such as InstructBLIP [9] and BLIP [16] primarily rely on high-quality paired language/image datasets (e.g. VQAv2 [12], VizWiz [13], TextCaps [28], etc.). However, these datasets are expensive to collect and challenging to adapt for broader coverage due to the need for manual text annotation. On the other hand, initiatives such as LLaVA [20] train on perception and simple caption datasets in conjunction with the reasoning capabilities of LLMs to semi-automatically generate synthetic conversational ground truth. Unfortunately, such outputs can also contain nonfactual statements suffering from hallucinations, omissions, and inaccuracies in attribute or relational descriptions, as they are generated by text-only models based on sparse information about the actual image. Hence, the resulting trained model is still not ideal (see LLaVA’s output in Figure 1).
@@
Instead, we propose a novel and generally applicable framework using fine-grained reward modeling. It efficiently and substantially enhances the visual grounding of LVLMs beyond pre-trained baselines such as LLaVA, while simultaneously preserving their capability to generate extended and detailed descriptions. Given a pre-trained LVLM (e.g., LLaVA), we input a set of images with prompts and sample multiple text outputs. Human annotators evaluate each image-text pair and assign fine-grained, per-sentence evaluation scores based on the quality of the generated text. This process results in a newly compiled dataset comprising image-text-evaluation trios. We train a reward model with this dataset and use it to fine-tune the pretrained LVLM (for further details, see Section  3). Notably, this fine-tuning process markedly improves the model’s visual grounding capabilities with just 16K data samples, demonstrating the method’s efficiency.
@@
To further improve the performance of our system at negligible cost, we also develop an automatic method to build the reward model without additional human effort, which proves to be highly effective in improving the visual grounding capabilities of LVLMs.
@@
Finally, we amalgamated the strengths of both reward models to develop a complete solution, which we refer to as ViGoR (Visual Grounding Through Fine-Grained Reward Modeling) throughout the remainder of the paper. To assess ViGoR’s efficacy, we evaluate it against two commonly used benchmarks, POPE [18] and MME [11], where it demonstrated significant advancements over baseline models. Moreover, we construct a comprehensive and challenging dataset, MMViG, specifically designed to validate the visual grounding capabilities of large vision language models. Finally, we use the recently released GPT-4V to automatically assist in evaluating our method.
In summary, our contributions are fourfold.
@@
  • We introduce a novel framework that incorporates fine-grained reward modeling, substantially enhancing the visual grounding of LVLMs.
@@
  • We develop reward models that require little human effort while leveraging the impressive advances in powerful and robust visual perception, demonstrating marked improvements in visual grounding efficiency.
  • We construct a comprehensive and challenging dataset, MMViG, specifically designed to validate the visual grounding capabilities of large vision language models.
@@
  • We created and plan to release the human evaluation dataset comprising 16K pairs of images and generated results, along with fine-grained evaluation annotation providing detailed quality assessments.
@@ 2 Related Work
Large Vision Language Models
In the domain of large vision language models (LVLMs), recent advances have been remarkable. The integration of large language models (LLM) such as GPT [4, 24], PaLM [8], BLOOM [32], LLaMA [30], and Vicuna [7] has significantly advanced the field. Flamingo [1] and its open source counterpart OpenFlamingo [3], along with IDEFICS, have been pivotal in integrating LLMs with vision-language pretraining, employing innovative techniques like gated cross-attention dense blocks. PaLI’s [6] research on the scaling of vision and language (V&\&&L) components across various tasks has been instrumental. Additionally, PaLM-E’s [10] extension of LMM to the embodied domain, and BLIP-2’s [16] introduction of the Querying Transformer (Q-former) to align image and language encoders mark significant progress. InstructBLIP [9] further enhances this approach. Otter’s [15] focus on improving OpenFlamingo’s instruction-following capabilities and MiniGPT-4’s [35] recommendation for a single project layer to align visual and linguistic models demonstrate its efficiency and capability. mPLUG-Owl [33] introduces a novel approach by aligning visual characteristics before fine-tuning the language model using LoRA [14]. LLaVA directly injects visual tokens into a pre-trained LLM and finetunes the model with synthetic conversations generated by GPT-4 using metadata and short captions as input. While these techniques primarily focus on architectural innovation for aligning image features within the large language model domain or leveraging extensive image-text data for instruction tuning, our work introduces a distinct and novel general framework designed to enhance the visual grounding capability of any LVLM.
Visual Perception Models
Recent advances in visual perception models have demonstrated remarkable proficiency in handling a variety of tasks in open world scenes. In particular, CLIP [25] has exhibited a robust capability for zero-shot classification. By reformulating object detection as a grounding problem and incorporating additional grounding data, GLIP [17] has achieved semantic alignment at both the phrase and region levels, leading to impressive open-set detection performance. GroundingDINO [21] represents another significant step, grounding the state-of-the-art transformer-based object detector DINO [5] with language pretraining for open-set generalization. This model demonstrates a strong ability to discern whether elements are present in a scene and accurately detect object counts.
However, such advanced visual grounding abilities have not yet been fully realized in current LVLMs. To bridge this gap, we devised an automatic method for building a reward model using these vision perception models. Our approach aims to distill their sophisticated visual grounding capabilities directly into LVLMs.
Reward Modeling
@@
However, in the realm of Large Vision Language Models (LVLMs), the application of reward modeling remains underexplored, with most existing work focusing predominantly on instruction tuning [20, 35]. One exception is LLaVA-RLHF which adapts the Reinforcement Learning from Human Feedback (RLHF) from the text domain to the task of vision-language alignment, where human annotators are asked to compare two responses and pin-point the more hallucinated one [29]. Drawing inspiration from LLaMA-2, we propose the integration of reward modeling with a rejection sampling strategy into the LVLM training framework. Different from LLaVA-RLHF which uses a sparse reward model, our aim is to leverage fine-grained reward models to substantially improve the visual grounding capabilities of LVLMs, potentially leading to more accurate and contextually relevant output in vision-language tasks.
@@ 3 ViGoR: Visual Grounding Improvement Framework
Our primary goal is to increase visual grounding and reduce hallucinations while maintaining the strong intuitive reasoning and creative thought process of pre-trained LLMs and LVLMs. While it has been shown that high quality human annotations for supervised fine-tuning is the most straightforward approach for significantly improving the performance of LVLMs, it is cost-prohibitive for many application scenarios. As such, we wish to efficiently leverage annotator time and the latest advances in direct visual perception models such as powerful open-set object detectors.
To achieve this, we construct a system to fine-tune a base LVLM with rejection sampling, similar to LLaMA2 [31], allowing the model to improve through using intelligent ranking of its own sampled output. This requires a robust and perceptive scoring system for the generated text. In our work, we use two complementary solutions in parallel. In particular, we train a reward model to incorporate human annotator assessments of text outputs from the LVLM, and provide positive and negative assessments of the LVLM during training time with unlabeled examples. As well, we leverage an open-set object detector and heuristics to verify the existence or absence in the image of the named noun entities extracted from the generated descriptions. Finally, we combine these signals into a single reward score, and use it to select the best description among the initial samples. This sample undergoes additional refinement and is used for the supervised fine-tuning of the LVLM. We refer the reader to Figure 2 for a visualization.
@@ 3.1 Reward Modeling via Fine-Grained Human Feedback
Using the collected annotations, we fine-tune a dedicated LVLM as the reward model on the annotations using instruction tuning to judge the base LVLM’s generated results during training time. The reward model is trained to output a string which encodes the various scores given the underlying image and the LVLM’s output as the input. Unlike existing work, we train the model to provide sentence-level human evaluations. To provide the necessary reference context for scoring each sentence, we prepend the sentence with all preceding generated text and explicitly ask the reward model to score the last sentence in the given passage. For example, a typical prompt is “Assess the accuracy of the last sentence in the following description of this image, and return a single number.” Due to the fine-grained feedback provided by the annotators, these prompts result in either a positive response (when no errors are found in the sentence), or a detailed negative response (where one of several error types are found in the sentence).
As we show in the ablation studies, this method significantly improves the reward model’s capabilities to guide the fine-tuning of the LVLM. Furthermore, we see that the reward model trained with fine-grained feedback can better understand the link between errors in the descriptions and the image, resulting in superior fine-tuning performance.
@@ 3.2 Reward Modeling with Automatic Methods
Beginning with a caption generated by the LVLM from an image, we first identify the individual nouns mentioned using standard named noun entity recognition. Next, we employ an open-set object detector, prompted with these nouns and the input image, to verify the existence of the objects. A correct identification is rewarded with a positive score, while hallucinated objects incur a penalty. For this purpose, we use Grounding DINO [21] for its strong performance across a wide variety of image domains. However, we note that while these detection models are adept at identifying the existence of objects, their ability to detect other types of errors is limited (e.g., object attribution, object relationship…). This limitation underscores the continued necessity for human-preference-based reward modeling.
@@ 3.3 Reward Score and Rejection Sampling
As both our reward model-based and automated methods provide fine-grained reward scores, we must design a strategy to combine the scores into a single score to enable rejection sampling at the description level. Note that this process embodies the same ambiguity faced by annotators in existing work, where they are asked to analyze complete text output in detail and rank them based on holistic preference. However, our solution combines the detailed analysis in a more consistent fashion.
We iterate through the samples and find the best sample to use as ground truth for supervised fine-tuning. As our primary focus is to reduce generation of erroneous descriptions, we use the signals that indicate errors as the primary selector. For each sample, we aggregate all negative signals from the two streams of description evaluation by normalizing their values with their respective variances and linearly combine them. We select the sample with the best penalty score as the best candidate, and use the positive scores as a tiebreaker when necessary.
@@ 3.4 Refinement Module
After selecting the best candidate from the rejection sampling process, we design a simple refinement module to further polish this selected candidate description to use as a regression target.
For each sentence within the description, we assess the presence of noun phrases deemed nonexistent by the reward modeling module. Should a sentence contain any such noun phrases, we eliminate that sentence entirely from the description. For instance, consider the description:“The image features a black Dodge Charger parked on a brick road next to a sidewalk. A person is walking in front of the car.” If the reward modeling module identifies “person” as a non-existent noun phrase, the sentence “A person is walking in front of the car.” is removed.
This refinement process, as evidenced in Table 5, significantly improves the visual grounding capability of the model. It effectively reduces the hallucination problem by removing inaccurate elements, thereby leading to notable improvements in the model’s performance.
@@ 3.5 Model Training
The resulting description from the refinement module is used as ground truth in supervised fine-tuning using the standard autoregressive objective as in the original LLaVA [20]. As will be demonstrated by the ablation studies, two signal sources are complementary and provide better results than either one alone. This overall process is visualized in Figure 2.

### v3

@@ Abstract
By combining natural language understanding, generation capabilities, and breadth of knowledge of large language models with image perception, recent large vision language models (LVLMs) have shown unprecedented visual reasoning capabilities. However, the generated text often suffers from inaccurate grounding in the visual input, resulting in errors such as hallucination of nonexistent scene elements, missing significant parts of the scene, and inferring incorrect attributes of and relationships between objects. To address these issues, we introduce a novel framework, ViGoR (Visual Grounding Through Fine-Grained Reward Modeling) that utilizes fine-grained reward modeling to significantly enhance the visual grounding of LVLMs over pre-trained baselines. This improvement is efficiently achieved using much cheaper human evaluations instead of full supervisions, as well as automated methods. We show the effectiveness of our approach through a variety of evaluation methods and benchmarks. Additionally, we released our human annotation (https://github.com/amazon-science/vigor) comprising 15,440 images and generated text pairs with fine-grained evaluations to contribute to related research in the community.
@@ 1 Introduction
Figure 1: An illustration of inaccurate visual grounding in Large Vision Language Models. BLIP-2 [18] failed to provide detailed image descriptions. LLaVA [22]’s more detailed description contains both correct sentences as well as sentences with hallucinations and inaccurate inference. In contrast, our model preserves the logical reasoning and creativity of LVLMs while exhibiting significantly better accuracy and detail.
@@
Large language models (LLMs) have garnered intense interest throughout the research and academic communities. Typically, these models are pre-trained with tremendous amounts of automatically aggregated text data followed by further fine-tuning with specific user examples or evaluation feedback. This process enables the models to follow human-provided prompts and retrieve useful relevant information or solve logical problems. Recently, numerous techniques [18, 10, 38, 22, 34, 17, 35] have enhanced these breakthroughs with the ability to understand visual information by further integrating image features into the prompt encoding process. Although these works have successfully aligned image features into the large language model domain, they still exhibit significant problems. While a strong contextual grounding of language-only models can be learned from the enormous corpus of text data, paired training data for multimodal language/vision models is more limited, while the complexity of the task is arguably higher as the model must align two disparate modalities. Large, automatically compiled datasets, such as the LAION-5B dataset [28], tend to feature only simple images with very short text descriptions. Training with such data often leads large vision language models to fail in capturing the essential details of the image, returning only a short and coarse description (see BLIP-2 [18] output in Figure 1). Moreover, techniques such as InstructBLIP [10] and BLIP [18] primarily rely on high-quality paired language/image datasets (e.g. VQAv2 [13], VizWiz [14], TextCaps [29], etc.). However, these datasets are expensive to collect and challenging to adapt for broader coverage due to the need for manual text annotation. On the other hand, initiatives such as LLaVA [22] train on perception and simple caption datasets in conjunction with the reasoning capabilities of LLMs to semi-automatically generate synthetic conversational ground truth. Unfortunately, such outputs can also contain non-factual statements suffering from hallucinations, omissions, and inaccuracies in attribute or relational descriptions, as they are generated by text-only models based on sparse information about the actual image. Hence, the resulting trained model is still not ideal (see LLaVA’s output in Figure 1).
@@
Instead, we propose a novel and generally applicable framework using fine-grained reward modeling. It efficiently and substantially enhances the visual grounding of LVLMs beyond pre-trained baselines such as LLaVA, while preserving their capability to generate extended and detailed descriptions. Given a pre-trained LVLM (e.g., LLaVA), we input a set of images with prompts and generate multiple text outputs. Human annotators are asked to evaluate each image-text pair. As seen in the LLaVA output in Figure 1, a lengthy description generated by a competent LVLM can contain both correct and incorrect sentences. Attempting to assign a single holistic score is ambiguous for the annotator, as well as a subsequently trained reward model. In our design, annotators assign fine-grained, per-sentence evaluation scores. This results in a newly compiled dataset comprising image-text-evaluation trios. We train a reward model to also predict dense reward scores and use it to fine-tune the pre-trained LVLM (see Section  3). This fine-tuning process markedly improves the model’s visual grounding capabilities with just 16K data samples, demonstrating the method’s efficiency.
@@
To further improve the performance of our system at negligible cost, we also develop a reward model scheme based on automatic methods without additional human effort, which is proven to be highly effective in improving the visual grounding capabilities of LVLMs.
@@
Finally, we amalgamated the strengths of both reward models to develop a complete solution, which we refer to as ViGoR (Visual Grounding Through Fine-Grained Reward Modeling) throughout the remainder of the paper. To assess the efficacy of ViGoR, we evaluate it against two commonly used benchmarks, POPE [20] and MME [12], where it demonstrated significant improvements over the baseline models. Finally, we compare the performance benefit of our approach to baselines for a variety of tasks that require accurate visual grounding. In summary, we make the following three major contributions.
@@
  • We introduce a novel framework that incorporates fine-grained reward modeling with easily implemented rejection sampling, substantially enhancing the visual grounding of LVLMs.
@@
  • We develop reward models that require little human effort while leveraging the impressive advances in powerful and robust visual perception models, demonstrating marked improvements in visual grounding efficiency.
@@
  • We create and release the human evaluation dataset comprising 15.4K pairs of images and generated results, as well as the fine-grained assessments of the latter by our annotators.
@@ 2 Related Work
Large Vision Language Models.
Recent advances in LVLMs have been remarkable, such as the integration of large language models (LLM) such as GPT [5, 27], PaLM [9], BLOOM [33], LLaMA [31], and Vicuna [8]. Flamingo [1] and its open source counterpart OpenFlamingo [3], along with IDEFICS, have been pivotal in integrating LLMs with vision-language pretraining, using techniques like gated cross-attention dense blocks. PaLI’s [7] research on the scaling of vision and language (V&\&&L) components across various tasks has been instrumental. As well, PaLM-E’s [11] extension of LLM to the embodied domain, and BLIP-2’s [18] introduction of the Querying Transformer (Q-former) to align image and language encoders mark significant progress. InstructBLIP [10] further enhances this approach. Otter’s [16] improves OpenFlamingo’s instruction-following capabilities and MiniGPT-4’s [38] recommendation for a single project layer to align visual and linguistic models demonstrate its efficiency and capability. mPLUG-Owl [34] aligns visual characteristics before fine-tuning the language model using LoRA [15]. LLaVA directly injects visual tokens into a pre-trained LLM and finetunes the model with synthetic conversations generated by GPT-4 using metadata and short captions as input. While these techniques primarily focus on architectural innovation for aligning image features to the feature space of LLMs or leveraging extensive image-text data for instruction tuning, our work introduces a distinct and novel general framework designed to enhance the visual grounding capability of any LVLM.
Visual Perception Models.
Recent advances in visual perception models have demonstrated remarkable proficiency in handling a variety of tasks in open world scenes. In particular, CLIP [26] has exhibited a robust capability for zero-shot classification. By reformulating object detection as a grounding problem, GLIP [19] has achieved semantic alignment at both the phrase and region levels, leading to impressive open-set detection performance. GroundingDINO [23] represents another significant step, grounding the state-of-the-art transformer-based object detector DINO [6] with language pretraining for open-set generalization. This model demonstrates a strong ability to discern whether elements are present in a scene and accurately detect object counts.
However, such advanced visual grounding abilities have not yet been fully realized in current LVLMs. To bridge this gap, we devise an automatic method for building a reward model using these vision perception models, and distill their strong visual grounding capabilities directly into LVLMs.
Reward Modeling.
@@
However, in the realm of Large Vision Language Models (LVLMs), the application of reward modeling remains underexplored, with most existing work focusing predominantly on instruction tuning [22, 38]. One exception is LLaVA-RLHF which adapts the Reinforcement Learning from Human Feedback (RLHF) from the text domain to the task of vision-language alignment, where human annotators are asked to compare two responses and pin-point the more hallucinated one [30]. Inspired by LLaMA-2, we combine reward modeling with rejection sampling in the LVLM training framework. While LLaVA-RLHF uses a reward model that produces sparse signals, we leverage fine-grained reward models to improve the visual grounding capabilities of LVLMs, leading to more accurate and contextually relevant output in vision-language tasks.
@@ 3 ViGoR: Visual Grounding Improvement Framework
Our primary goal is to increase visual grounding and reduce hallucinations while keeping the strong intuitive reasoning and creative thought process of pre-trained LLMs and LVLMs. While it has been shown that high quality human annotations for supervised fine-tuning is a straightforward approach for significantly improving LVLMs, it is cost-prohibitive for many application scenarios. As such, we wish to efficiently leverage annotator time and the latest advances in direct visual perception models such as powerful open-set object detectors.
We construct a system to fine-tune a base LVLM with rejection sampling, similar to LLaMA-2 [32], allowing the model to improve through using intelligent ranking of its own sampled output. This requires a robust and perceptive scoring system for the generated text. In our work, we use two complementary solutions in parallel. In particular, we train a reward model to incorporate human annotator assessments of text outputs from the LVLM, and provide positive and negative assessments of the LVLM during training time with unlabeled examples. As well, we leverage an open-set object detector and heuristics to verify the existence or absence in the image of the named noun entities extracted from the generated descriptions. Finally, we combine these signals into a single reward score, and use it to select the best description among the initial samples. This sample undergoes additional refinement and is used for the supervised fine-tuning of the LVLM. We refer the reader to Figure 2 for a visualization.
@@ 3.1 Reward Modeling via Fine-Grained Human Feedback
Using the collected annotations, we fine-tune a dedicated LVLM as the reward model on the annotations using instruction tuning to judge the base LVLM’s generated results during training time. The reward model is trained to output a sequence of text tokens which encodes the various scores given the underlying image and the LVLM’s output as the input. While existing work [30] generates a single holistic score for the entire text output, this process can be ambiguous when the description contains both correct and incorrect components (see Figure 1). Instead, we train the model to produce sentence-level evaluations. This fine-grained approach reduces ambiguity, and increases the detailed visual grounding of the reward model. To provide the necessary context for scoring each sentence, we prepend the sentence with all preceding generated text and explicitly ask the reward model to score the last sentence (which is also the target sentence) in the given passage. A typical prompt is “Assess the accuracy of the last sentence in the following description of this image, and return a single number." Due to the fine-grained feedback provided by the annotators, these prompts result in either a positive response (when no errors are found in the sentence), or a detailed negative response (where one of several error types are found in the sentence). More details are found in the Supplementary Materials.
As we show in the ablation studies, compared with holistic-based method, this fine-grained method significantly improves the reward model’s capabilities to guide the fine-tuning of the LVLM. Furthermore, we see that the reward model trained with fine-grained feedback can better understand the link between errors in the descriptions and the image, resulting in superior fine-tuning performance.
@@ 3.2 Reward Modeling with Automatic Methods
Starting with a caption generated by the LVLM from an image, we identify the individual nouns mentioned using standard named entity recognition with NLTK [4]. Next, we prompt an open-set object detector with these nouns to verify the existence of the objects in the image. A correct identification is rewarded with a positive score, while hallucinated objects incur a penalty. We use Grounding DINO [23] for its strong performance across a wide variety of image domains. However, we note that while these detection models are adept at identifying the existence of objects, their ability to detect other types of errors is limited (e.g. object attributes and relationships). This limitation underscores the continued necessity for human-preference-based reward modeling.
@@ 3.3 Reward Score and Rejection Sampling
As both our reward model-based and automated methods provide fine-grained reward scores, we must design a strategy to combine them into a single score to enable rejection sampling at the description level. Note that this process embodies the same ambiguity faced by annotators in existing work, where they are asked to analyze complete text output in detail and rank them based on holistic preference. However, our solution combines the detailed analysis in a more principled and consistent fashion.
The sample with the best score is used as ground truth for supervised fine-tuning. As our focus is to reduce generation of erroneous descriptions, we use the signals that indicate errors as the primary selector. For each sample, we aggregate all negative signals from the two streams of description evaluation by normalizing their values with their respective variances and linearly combine them. We select the sample with the smallest penalty score as the best candidate, and use the positive scores as a tiebreaker when necessary.
@@ 3.4 Refinement Module
We design a simple refinement module to further polish the selected candidate description to use as a regression target (see part c of Figure 2). For each sentence within the description, we assess the presence of noun phrases deemed nonexistent by the reward modeling module. Should a sentence contain any such noun phrases, we eliminate that sentence entirely from the description. For instance, consider the description:“The image features a black Dodge Charger parked on a brick road next to a sidewalk. A person is walking in front of the car." If the reward modeling module identifies “person" as a non-existent noun phrase, the sentence “A person is walking in front of the car." is removed. This refinement process effectively reduces the hallucination problem by removing inaccurate elements, thereby leading to notable improvements in the model’s performance.
@@ 3.5 Model Training
The resulting description from the refinement module is used as ground truth in supervised fine-tuning with the standard autoregressive objective as in the original LLaVA [22]. As will be demonstrated by the ablation studies, two signal sources are complementary and provide better results than either one alone. This overall process is visualized in Figure 2.